OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

ok im really not sure since i did it last year but i think that u need to find common ratio. 781,250=10*r^7. take the 7th root so r=5. from there use ur geometric formula. 10(1-5^8) take ans divide by -4. ans=976560

OpenStudy (anonymous):

goemtric formula summation= 1st term * 1-common ratio^nth. take ans divide by 1-common ratio.